# A Survey of Graph Unlearning

### v3

@@ 5.5 Ethical Consideration in Transportation Networks
Commonly, machine learning models are employed in traffic analysis to predict traffic patterns, optimize routes, and enhance transportation systems [54]. However, these models may accidentally acquire sensitive information about an individual’s mobility patterns, daily commute routes, and specific location histories, thereby compromising privacy. The risk of data leakage and identification of individuals through these models raises ethical concerns regarding the protection of the anonymity and privacy of travelers’ data. By incorporating graph unlearning into traffic analysis, it is possible to effectively protect individual privacy. Using graph unlearning techniques permits the removal of sensitive data from trained models while preserving the models’ capacity to provide accurate traffic predictions and enhance transportation efficiency.
@@ 5.6 Ethical Consideration in Biological Networks
5.7 Individuals’ Privacy in Data Sharing in Collaborative Environment
@@ 5.8 Adversarial Setting
In sensitive domains like healthcare, finance, and personal recommendation systems, the consequences of a security breach or an attacker gaining access to a trained machine learning model can be severe. In these scenarios, an attacker may attempt to compromise user privacy by injecting malicious data into the model’s training dataset or by stealing sensitive information from the model itself. This malicious activity can lead to various detrimental outcomes, such as unauthorized access to personal medical records, financial data, or private user preferences [56]. For instance, in the healthcare domain, if an attacker gains access to a medical diagnosis model, they could inject misleading or harmful data, potentially leading to inaccurate diagnoses and jeopardizing patient safety. In the finance sector, unauthorized access to a fraud detection model could enable attackers to manipulate the system, bypass security measures, and conduct fraudulent transactions. Similarly, in personal recommendation systems, adversaries may exploit information about users’ preferences and behaviors obtained through malicious data injection or information stealing [57]. This could result in biased or manipulated recommendations, eroding user trust and leading to potential misuse of personal data. To safeguard against these risks, the application of machine unlearning becomes essential.
@@ 5.9 Limited Resource Environment
The application of graph unlearning may holds immense potential, especially in limited resource environments such as the Internet of Things (IoT). In such scenarios, devices at the edge may encounter situations where certain training examples become obsolete, out of distribution, or no longer valid [14, 58]. For instances, a network of smart traffic cameras placed strategically across the city to monitor traffic flow and detect any congestion or accidents. Each traffic camera acts as a node in the knowledge graph, and the connections between these nodes represent the spatial relationships between different camera locations. Assume that we have different devices at different locations of the cities to collect data from the nearby cameras. As vehicles move through the city, the smart traffic cameras continuously capture real-time data, such as vehicle count, speed, and license plate recognition. This data is then used to generate and update the knowledge graph, which dynamically represents the traffic patterns and conditions in the city. Over time, the traffic patterns may change due to various factors, such as construction work, events, or changes in city planning. Some of the historical data captured by the traffic cameras on some devices may become less relevant for predicting and managing the current traffic conditions. To optimize the use of resources and improve the accuracy of traffic predictions, graph unlearning can be applied to the knowledge graph. Here, the concept of fine-tuning from an unlearning perspective becomes valuable, particularly when the devices aim to retain their existing data without adding more. Instead, these devices may selectively flag certain types of data for unlearning, allowing them to adapt to dynamic changes in their operating environment. The idea of personalized or customized unlearning strategies tailored to each individual device’s needs further accentuates the versatility of graph unlearning.
@@ 6.1 Differential Privacy
The potential leakage of sensitive information in scenarios where the trained model is shared or used in sensitive applications raises concerns about individual privacy [60]. Differential Privacy (DP) offers solutions to this issue by introducing controlled noise into the model updates during SGD, ensuring that individual data points do not unduly influence the model’s parameters. Well-known works in this line of research include [61, 62], which introduces a novel algorithmic technique for training models under the umbrella of differential privacy with tight privacy bounds. [61] approach involves clipping the $l_{2}$ norm of each gradient, computing the average, adding noise to protect privacy, and then taking a step in the opposite direction of this average noisy gradient. Additionally, it incorporates a privacy loss function to further enhance privacy protection during model training.
In recent years, there have been several efforts to incorporate DP to provide formal privacy guarantees in various GNN learning scenarios. One notable proposal by [57] introduces an edge-level DP learning algorithm for GNNs. They achieve this by directly perturbing the input graph using randomized response (EdgeRand) or the Laplace mechanism (LapGraph) and subsequently train the GNN on the perturbed graph. Despite its effectiveness, this method does not readily extend to the node-level privacy setting. In a different approach, [63] present a locally private GNN model in a distributed learning setup, where node features and labels remain private while the GNN training is federated by a central server with access to graph edges. However, this method is not suitable for scenarios where graph edges are also private. In a similar vein, another approach [64], denoted as GAP, employs aggregation perturbation to achieve edge-level privacy. The GAP framework is composed of three key modules: the encoder module, which learns solely from node features, the aggregation module, responsible for learning node representations from the graph topology, and the classification module. Addressing node-level privacy in a centralized learning setting, [56] adapt the PATE framework [65]. They train the student GNN model using publicly labeled graph data and teacher GNN models trained exclusively for each query node. Nevertheless, their reliance on public graph data restricts the applicability of their approach.
A wave of recent approaches has emerged, aiming to tackle the issue of privacy through techniques like federated and split learning. [36] propose a GNN that leverages structural similarity and federated learning, striving to cloak content and structure information, thereby enhancing privacy protection. On a similar front, [37] crafted a distributed and secure framework, embarking on the journey to unravel object representations in video data through graph sequences, expertly blending GNN and federated learning. They wove in secure aggregation primitives to fortify privacy during the federated learning process. However, their focus lies in a different territory, as they assume each party possesses a series of graphs extracted from video data, and the server orchestrates federated learning to construct an inductive GNN over this distributed dataset of graphs. Meanwhile, [66] directed their efforts towards privacy-preserving node classification. They splintered the computation graph of a GNN across multiple data holders, enlisting the expertise of a trusted server to amalgamate information from diverse parties and achieve comprehensive training. Nonetheless, these approaches depend on a trusted third party for model aggregation, leaving their privacy guarantees without formal validation. Recognizing that privacy requirements may differ for node attributes and graph structure in graph unlearning, [67] introduce a graph deferential privacy (GDP)framework. This paper also introduces the concept of relaxed node-level data adjacency. This relaxation allows to establish varying degrees of privacy for the graph topology while still ensuring the privacy of individual node attributes. With this relaxation, GDP framework provides a flexible solution to address diverse privacy needs in graph-based learning scenarios. In a related study by [19] , an alternative method involving shard graphs within the context of differential privacy has been put forward. This proposal leverages lightweight cross-attention adapters, which operate on the representations of a transformer encoder that has been pretrained. Specifically, each node in the graph is linked to an InCA adapter, as introduced in [68], which is trained on the combined data of the node and all its outward edges.
As we observe, there has been a surge of approaches in the related domain. However, there is still a pressing need to advance this field in several new directions. We provide a few potential research directions and discuss them in the following.
@@ 6.2 Approximate Graph Unlearning for Non-Convex Settings
While remarkable progress has been made in graph unlearning, a notable area that warrants further exploration is the development of approximate unlearning methods tailored to non-convex settings. Currently, many existing approaches primarily focus on convex settings, where optimization landscapes are well-behaved and lend themselves to more straightforward solutions [20, 34]. However, real-world graph-based machine learning applications often involve non-convex settings with complex and irregular optimization landscapes. As such, designing effective approximate unlearning techniques for these scenarios is challenging but crucial. The need for approximate unlearning in non-convex settings stems from the realization that many practical graph-based models operate in environments where convexity assumptions do not hold. Examples include MPNN and other highly expressive models, which are capable of capturing intricate patterns and dependencies within the data [69, 70]. In these complex settings, unlearning approaches need to cope with non-convex optimization landscapes, which introduce additional intricacies and make the task significantly more demanding [2].
Tackling approximate unlearning in non-convex settings is essential for the wider adoption of graph-based machine learning in various real-world applications [9]. The ability to efficiently remove sensitive information and adjust models in response to evolving privacy requirements will instill confidence among users and stakeholders in utilizing graph-based machine learning solutions. While this direction poses challenges due to the inherent difficulty of approximating models in non-convex settings, it presents an exciting opportunity to push the boundaries of graph unlearning research [21]. Solutions that can effectively handle non-convexity while maintaining a balance between privacy preservation and model performance will be invaluable in applications ranging from social network analysis and recommendation systems to bioinformatics and beyond.
@@ 6.5 Leveraging Graph Descriptors for Graph-level Unlearning
Several graph descriptor methods have emerged in recent years, each with its unique characteristics. For further reading, we refer the reader to recent works on graph descriptors [74, 73, 72, 75, 71]. By incorporating graph descriptors into graph-level unlearning, researchers can take advantage of the rich representations they offer, allowing for efficient unlearning without the need to retrain the model repeatedly. The fixed nature of graph embeddings ensures that graph-level privacy-preserving operations do not interfere with the overall performance and generalization capabilities of the machine learning model. Exploring novel graph descriptor techniques tailored to graph-level unlearning requirements is an exciting avenue for future research. Additionally, comparative studies that evaluate different graph embedding methods in the context of unlearning tasks will provide valuable insights into their strengths and limitations. Moreover, since these descriptors are quite efficient, new embeddings could be quickly computed if node or edge-level removal requests are received.
@@ 6.7 Unlearning for Foundation Models
With the success of large language models (LLMs) [76], large-scale foundation models have been developed across various domains, including graphs [77, 78]. These models typically follow a common paradigm of pre-training followed by fine-tuning. However, despite advancements in efficient unlearning algorithms, directly unlearning from the pre-trained model remains time-consuming due to its large size. This highlights the need to explore unlearning during the fine-tuning stage. Furthermore, retrieval-augmented generation (RAG) has been employed to enhance the performance of foundation models, but RAG relies on external datasets that might contain data a user seeks to remove. Many studies leverage knowledge graphs as the external data source [79, 80] due to the powerful representation capabilities of graphs, making the challenges of graph unlearning unavoidable in these graph-based RAG systems. Therefore, investigating unlearning within the RAG component, particularly in graph-based systems, rather than focusing solely on the foundation model, presents an intriguing research direction.
@@ 7 Conclusion
This paper offers a comprehensive review of the emerging field of graph unlearning, which plays a vital role in responsible AI development. It details various graph unlearning approaches and methodologies, along with their applications in diverse domains, highlights the wide-ranging impact of this research area, including social networks, financial networks, and transportation systems. We have emphasized the sensitivity of GML to data privacy and adversarial attacks, underscoring the importance of graph unlearning techniques in addressing these critical concerns. Additionally, the connections drawn between graph unlearning and differential privacy underscore the significance of privacy-preserving techniques in the responsible design and deployment of AI systems. The comprehensive taxonomy, up-to-date literature overview, and evaluation measures provided in this survey empower researchers to explore new approaches and contribute to this evolving field. The suggested research directions offer exciting opportunities for future investigations, encouraging the development of novel algorithms, fortified privacy guarantees, and exploration of applications in emerging domains.

### v4

@@ 5.5 Ethical Consideration in Transportation Networks
Commonly, machine learning models are employed in traffic analysis to predict traffic patterns, optimize routes, and enhance transportation systems [72]. Such models may also capture sensitive details about individuals’ mobility patterns, commute routes, and location histories, creating privacy risks. Graph unlearning can remove this information from trained models, safeguarding traveler anonymity while maintaining the accuracy and efficiency of traffic predictions.
@@ 5.6 Ethical Consideration in Biological Networks
5.7 Individuals’ Data Privacy in Collaborative Environment
@@ 5.8 Adversarial Setting
In sensitive domains like healthcare, finance, and personal recommendation systems, the consequences of a security breach compromising access to a trained machine learning model can be severe. In these scenarios, an attacker may attempt to jeopardize user privacy by injecting malicious data into the model’s training dataset or by stealing sensitive information from the model itself. This malicious activity can lead to various detrimental outcomes, such as unauthorized access to personal medical records, financial data, or private user preferences [75]. For instance, in the healthcare domain, if an attacker gains access to a medical diagnosis model, they could inject misleading or harmful data, potentially leading to inaccurate diagnoses and jeopardizing patient safety. In the finance sector, unauthorized access to a fraud detection model could enable attackers to manipulate the system, bypass security measures, and conduct fraudulent transactions. Similarly, in personal recommendation systems, adversaries may exploit information about users’ preferences and behaviors obtained through malicious data injection or information stealing [76]. This could result in biased or manipulated recommendations, eroding user trust and leading to potential misuse of personal data. To safeguard against these risks, several recent methods have been proposed to mitigate the effects of adversarial attacks on GNNs. For instance, in [12], the authors have introduced a three-step approach. First, attack detection methods are employed to identify anomalous nodes or edges within the graph. The accessibility of these anomalous nodes or edges varies depending on the specific scenario and the level of awareness regarding poisoned samples. Second, a fine-tuned subgraph is constructed based on the detected anomalous nodes or edges, ensuring that the subgraph is optimized for subsequent processing. Finally, the constructed subgraph is utilized to adjust or re-optimize the parameters of the poisoned GNN model, aiming to restore its predictive accuracy while mitigating the influence of adversarial perturbations. In order to allow the utilization of new attack advancements, [77] recently presented a game-theoretic approach that integrates membership inference attack (MIAs) into the framework of unlearning algorithms. In this setting, the unlearning task is formulated as a Stackelberg game, where the unlearner aims to remove specific training data from a model, while the auditor leverages MIAs to detect any residual evidence of the supposedly erased data.
@@ 5.9 Limited Resource Environment
The application of graph unlearning may holds immense potential, especially in limited resource environments such as the Internet of Things (IoT). In such scenarios, devices at the edge may encounter situations where certain training examples become obsolete, out of distribution, or no longer valid [78, 79]. For instances, a network of smart traffic cameras placed strategically across the city to monitor traffic flow and detect any congestion or accidents. These cameras usually share the same on-device model, which can be synchronously finetuned with federated learning as new data arrives. Moreover, each traffic camera acts as a node in the knowledge graph, and the connections between these nodes represent the spatial relationships between different camera locations. Assume that we have different devices at different locations of the cities to collect data from the nearby cameras. As vehicles move through the city, the smart traffic cameras continuously capture real-time data, such as vehicle count, speed, and license plate recognition. This data is then used to generate and update the knowledge graph, which dynamically represents the traffic patterns and conditions in the city. Over time, the traffic patterns may change due to various factors, such as construction work, events, or changes in city planning. Some of the historical data captured by the traffic cameras on some devices may become less relevant for predicting and managing the current traffic conditions. These observations incentivize the development of a graph unlearning method that, when applied to the distributed knowledge graph formed by the cameras, can optimize the use of resources and improve the accuracy of traffic predictions. Here, the concept of fine-tuning from an unlearning perspective becomes valuable, particularly when the devices aim to retain their existing data without adding more. Instead, these devices may selectively flag certain types of data for unlearning, allowing them to adapt to dynamic changes in their operating environment. Furthermore, IoT devices have limited computational resources, which in turn gives a push to focus on unlearning efficiency. The common approach of periodically retraining all models become less feasible for the modern AI landscape, where large (foundation) models become more common, and edge devices such as smartphones become the norm. As a result, efficient graph unlearning algorithms would prove to be very beneficial as the field of ML keep progressing at a break neck pace.
@@ 6.1 Data Privacy
The potential leakage of sensitive information in scenarios where the trained model is shared or used in sensitive applications raises concerns about individual privacy [82]. Differential Privacy (DP) offers solutions to this issue by introducing controlled noise into the model updates during SGD, ensuring that individual data points do not unduly influence the model’s parameters. Well-known works in this line of research include [83, 84], which introduces a novel algorithmic technique for training models under the umbrella of DP with tight privacy bounds. [83] approach involves clipping the $l_{2}$ norm of each gradient, computing the average, adding noise to protect privacy, and then taking a step in the opposite direction of this average noisy gradient. Additionally, it incorporates a privacy loss function to further enhance privacy protection during training. Any DP approach can then be trivially applied to unlearning, as the former is a strictly stronger definition of the latter [85].
Despite its desired theoretical guarantee, DP is not suitable for many cases given it requires careful prior planning. Solving this issue, certified removal methods aims to provide the same guarantee for when training was conducted without DP. Beyond [53] that served as foundation for [34] to introduce certified unlearning on graphs, literature on this topic for general machine learning has moved forward while its graph-equivalent did not receive such interest. [86] similarly utilizes second-order updates to unlearn samples, but provides guarantee with respect to the test loss instead of the gradient norm. [87] extends the previous ideas to work with non-convexity by proposing an efficient inverse Hessian approximation for the update. Attacking from another angle, by embracing DP and noisily finetuning model on unforgotten data, [88] utilizes gradient clipping and model clipping to circumvent the loss smoothness requirement, in turn bounding sensitivity and injected noise magnitude. While certified removal is not as strong as exact unlearning, the certification could provide better credibility for weaker approximate unlearning algorithms through a theoretical guarantee; and thus exploring this underutilized area of research in the graph context may prove to be very valuable, especially in privacy-critical scenarios.
Naturally, developments in general ML privacy have been adapted to GNNs as they gain prominence. Recent efforts apply DP to provide formal guarantees in various GNN scenarios. For edge-level privacy, [76] perturbs the input graph using randomized response (EdgeRand) or the Laplace mechanism (LapGraph) before training, though the method does not extend easily to node-level privacy. [89] propose a locally private GNN in a federated setting where node features and labels remain private but edges are shared, limiting applicability when edges are sensitive. The GAP framework [90] uses aggregation perturbation to protect edges, while [75] adapt the PATE framework [91] for node-level privacy, though reliance on public data constrains its use.
Privacy has also been addressed via federated and split learning. [92] combine structural similarity with FL to obscure content and structure, and [93] develop a secure FL framework for video-derived graph sequences using secure aggregation. [94] split GNN computation across data holders with a trusted server, but dependence on such a party limits guarantees. Recognizing differing privacy needs for node attributes and structure, [95] propose the GDP framework, introducing relaxed node-level data adjacency to balance topology and attribute privacy. [13] present an alternative using shard graphs with lightweight cross-attention adapters [96], linking each node to an adapter trained on its own and outward-edge data.
@@ 6.2 Approximate Graph Unlearning for Non-Convex Settings
While remarkable progress has been made in graph unlearning, a notable area that warrants further exploration is the development of approximate unlearning methods tailored to non-convex settings. Currently, many existing approaches primarily focus on convex settings, where optimization landscapes are well-behaved and lend themselves to more straightforward solutions [14, 38]. However, real-world graph-based machine learning applications often involve non-convex settings with complex and irregular optimization landscapes. As such, designing effective approximate unlearning techniques for these scenarios is challenging but crucial. The need for approximate unlearning in non-convex settings stems from the realization that many practical graph-based models operate in environments where convexity assumptions do not hold. Examples include MPNN and other highly expressive models, which are capable of capturing intricate patterns and dependencies within the data [97, 98]. In these settings, unlearning approaches need to cope with non-convex landscapes, which introduce additional intricacies making the task significantly more demanding [2].
Tackling approximate unlearning in non-convex settings is essential for the wider adoption of graph-based machine learning in various real-world applications [99]. The ability to efficiently remove sensitive information and adjust models in response to evolving privacy requirements will instill confidence among users and stakeholders in utilizing graph-based machine learning solutions. While this direction poses challenges due to the inherent difficulty of approximating models in non-convex settings, it presents an exciting opportunity to push the boundaries of graph unlearning research [16]. Solutions that can effectively handle non-convexity while maintaining a balance between privacy preservation and model performance will be invaluable in applications ranging from social network analysis, recommendation systems, to bioinformatics and beyond.
@@ 6.5 Leveraging Graph Descriptors for Graph-level Unlearning
Several graph descriptor methods have emerged in recent years, each with its unique characteristics. For further reading, we refer the reader to recent works on graph descriptors [105, 104, 102]. By incorporating graph descriptors into graph-level unlearning, researchers can take advantage of the rich representations they offer, allowing for efficient unlearning without the need to retrain the model repeatedly. The fixed nature of graph embeddings ensures that graph-level privacy-preserving operations do not interfere with the overall performance and generalization capabilities of the machine learning model. Exploring novel graph descriptor techniques tailored to graph-level unlearning requirements is an exciting avenue for future research. Additionally, comparative studies that evaluate different graph embedding methods in the context of unlearning tasks will provide valuable insights into their strengths and limitations. Moreover, since these descriptors are quite efficient, new embeddings could be quickly computed if node or edge-level removal requests are received.
@@ 6.7 Unlearning for Foundation Models
With the success of large language models (LLMs) [106], large-scale foundation models have been developed across various domains, including graphs [107, 108]. These models typically follow a common paradigm of pre-training followed by fine-tuning. However, despite advancements in efficient unlearning algorithms, directly unlearning from the pre-trained model remains time-consuming due to its large size. This highlights the need to explore unlearning during the fine-tuning stage. Retrieval-augmented generation (RAG) has been commonly employed to enhance the performance of foundation models, but RAG relies on external datasets that might contain data a user seeks to remove, as seen in Figure 3. Many studies leverage knowledge graphs as the external data source [109, 110] due to the powerful representation capabilities of graphs, making the challenges of graph unlearning unavoidable in these graph-based RAG systems. More recently, a new LLM-inspired approach to finetune foundation graph model has been graph prompt tuning [111, 112], where most of the model is frozen with the exception of a small subset of weights that directly modify the input graph itself—these weights are then updated to work with the target downstream task. This naturally creates the demand for adapting existing unlearning methods for LLMs [113, 114] to the graph setting, and analyzing whether existing pitfalls for LLMs [115] persist to the graph foundation models, presents an intriguing research direction.
@@ 7 Conclusion
This paper offers a comprehensive review of the emerging field of graph unlearning, which plays a vital role in responsible AI development. It details various graph unlearning approaches and methodologies, along with their applications in diverse domains, highlights the wide-ranging impact of this research area, including social networks, financial networks, and transportation systems. We emphasize the sensitivity of GML to data privacy and adversarial attacks, underscoring the importance of graph unlearning techniques in addressing these critical concerns. The comprehensive taxonomy, up-to-date literature overview, and evaluation measures provided in this survey empower researchers to explore new approaches and contribute to this evolving field. The suggested research directions offer exciting opportunities for future investigations, encouraging the development of novel algorithms, fortified privacy guarantees, and exploration of applications in emerging domains.